# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ 3.1 Introduction to TQFTs
A basic physical operation is the fusion of two anyons. We denote the fusion of anyons $a$ and $b$ as $a\times b$. In general, $a\times b$ is a linear combination of anyons. If $a\times b$ is a single anyon, for all $b$, we refer to $a$ as an Abelian anyon. It is clear that the set of all Abelian anyons $\cal A$ forms an Abelian group whose multiplication is given by fusion. The $n$-th power of an Abelian anyon $a^{n}$ is then defined in the obvious way.
@@
Let us return to the full set of anyons, not necessarily in $\cal A$. In a bosonic system, we require that $B$ be non-degenerate, i.e., the only anyon that braids trivially with all others is the identity $1$. (This requirement leads to the fact that the theory is modular. Hence, the M in MTC.) A fermionic theory is not modular. It includes a unique anyon $c$, besides $1$, that braids trivially with all the other anyons. It is the physical fermion, which satisfies $c^{2}=1$ and $h(c)=\frac{1}{2}$. Often we identify $c$ with $1$, since both are local particles.
Let us demonstrate the general discussion with the special case of Abelian theories. They are described by a $K$-matrix Chern-Simons Lagrangian [^7] There is a procedure for mapping the algebraic data of the TQFT to a $K$-matrix, which is given in App. C of [62]. We will not review it here, except to mention that it is possible to go back and forth.
$\begin{split}&\mathcal{L}={1\over 4\pi}\sum_{ij}K^{ij}b_{i}\wedge db_{j}\\ &K^{ij}\in\mathbb{Z}\,,\end{split}$ (3.3)
@@
Let us turn to the more general case. The topological spin $h(a)$ and the $\mathrm{U}(1)$ charge $Q(a)$ are subject to identifications due to attachment of local excitations. We consider three cases.
@@
    Physically, the ambiguity in the charge arises due to attaching a transparent boson with integer charge.
@@
    In a spin theory where $c$ is identified with the vacuum, [^9] A spin TQFT is a TQFT that is well-defined only on spin manifolds. And then, it depends on the choice of spin structure on that manifold [65, 66, 67]. Such a system does not have to be enriched by any global symmetry.
@@
    We prove in App. C that the different ways of assigning $\mathrm{U}(1)$ charge in this theory are essentially identical to the bosonic theory. They also correspond to different choices of Abelian vison, modulo fusion with $c$.
@@
    If $c$ is identified with the vacuum, then there is an additional equivalence relation which relates the tuple $(h(a),Q(a))$ to $(h(a\times c),Q(a\times c))$:
@@
    Just as $\hat{h}$, the $\pi$-flux defect is not just a convenient tool, but plays an important role in the physics. Classifying the different ways of assigning $\mathrm{U}(1)$ charges in a spinc TQFT can be done by extending the theory to a bosonic TQFT. Here, the $\pi$-flux line appears not as a defect line, but as an actual anyon in the theory. Once the theory has been extended to a bosonic TQFT, the techniques of [57] can be employed to classify charge assignments. We give much more detail in App. C.
@@ 3.2 Global symmetries, anomalies, and gauging
We now rephrase the discussion using the language of higher-form symmetries [37], within the framework of TQFT. The advantage of doing so is that the approach streamlines the matching of symmetries and anomalies between UV and IR theories, and the reasoning can be easily extended to non-topological theories with such symmetries.
@@
with $\lambda$ a flat $\mathrm{U}(1)$ gauge field, i.e., $d\lambda=0$, with $e^{is\oint\lambda}=1$ around every one-cycle. Such $\lambda$ is a $\mathbb{Z}_{s}$ gauge field and the global symmetry transformation (3.16) is a $\mathbb{Z}_{s}^{(1)}$ one-form global symmetry. [^12] Throughout this note, we use the convention that $\mathbb{Z}_{s}^{(p)}$ denotes a $\mathbb{Z}_{s}$ $p$-form symmetry.
Given a global symmetry, it is common to couple it to a background gauge field. In this case, the background gauge field $\cal B$ is a $\mathbb{Z}_{s}$ two-form field. This means that for every two-cycle $\cal C$ in our spacetime, $\int_{\cal C}{\cal B}$ is an integer modulo $s$. The effect of nonzero $\cal B$ is to twist the dynamical gauge field $b$. Instead of having $\int_{\cal C}{db\over 2\pi}\in\mathbb{Z}$, we have
@@
In other words, for a nonzero background field $\cal B$, the flux of $b$ has a fractional value.
@@
Regardless of our choice of ${M^{(4)}}$, with nonzero $\cal B$ the definition of the action depends on $b$ in the bulk ${M^{(4)}}$. Importantly, this is not as bad as it could have been because the dependence on $b$ in the bulk depends only on $\cal B$. This means that we can extend $b$ to the bulk in an arbitrary way, as long as we keep (3.17). Then, the dependence on the bulk is only through $\cal B$ and is given by a quadratic expression in $\cal B$, which we will discuss below. The significance of this statement is that even for nonzero $\cal B$, we can still view $b$ as a 2+1d field and the theory remains 2+1-dimensional.
What we have seen here is that the theory has a global symmetry acting as (3.16) and when we couple it to background gauge field $\cal B$, we need to specify additional information, which is captured by an extension of $\cal B$ to one dimension higher. This is the standard setting of an ’t Hooft anomaly. We conclude that the one-form $\mathbb{Z}_{s}^{(1)}$ global symmetry has an anomaly, which is described by [69, 37, 70, 68]:
@@
This discussion of $\cal B$ is related to the coupling to a background $\mathrm{U}(1)$ gauge field $A$ as in (3.15). To see that, note that in the absence of sources, (3.15) leads to the equation of motion of $b$
$sdb=dA\,$ (3.21)
@@
Hence, we identify $\int{dA\over 2\pi}=\int{\cal B}$, where $\cal B$ is the background field of the $\mathbb{Z}_{s}^{(1)}$ one-form global symmetry. For values of $\cal B$ satisfying this relation, the anomaly (3.18) is
@@
Let us try to gauge the one-form symmetry. The anomaly prevents us from doing it. However, if there are $m\in 2\mathbb{Z}$ and $n\in\mathbb{Z}$ such that $s=mn^{2}$, then, we can gauge $\mathbb{Z}_{n}^{(1)}\subset\mathbb{Z}_{s}^{(1)}$, because the anomaly for this subgroup vanishes. Denote the gauge field for this subgroup as $\mathfrak{b}$, in terms of which ${\cal B}={s\over n}\mathfrak{b}$ [^14] We use upper-case characters for classical background fields and lower-case characters for dynamical fields. Clearly, $\mathfrak{b}$ is a $\mathbb{Z}_{n}$ gauge field. Then, for fixed $\mathfrak{b}$, (3.17) becomes
@@
For example, the $\mathrm{U}(1)_{8}$ TQFT has a $\mathbb{Z}_{8}^{(1)}$ one-form global symmetry. A $\mathbb{Z}_{2}^{(1)}$ subgroup of it is anomaly-free and can be gauged. In the notation above, this corresponds to $s=8$ with $m=n=2$. This gauging leads to the $\mathrm{U}(1)_{2}$ TQFT.
We now generalize the gauging procedure and continue to focus on bosonic theories. A subset of the anyons in the TQFT $\cal T$ is the Abelian anyons ${\cal A}\subset{\cal T}$. They are the symmetry lines of the one-form global symmetry of $\cal T$. We consider an anomaly-free subset of them ${\cal A}_{0}\subset{\cal A}$. This means that for all $x,y\in{\cal A}_{0}$,
@@
Gauging ${\cal A}_{0}$ leads to another TQFT denoted by ${\cal T}/{\cal A}_{0}$. To find it, we couple $\cal T$ to background gauge fields for ${\cal A}_{0}$. If the symmetry associated with ${\cal A}_{0}$ is $\otimes_{I}\mathbb{Z}_{n_{I}}^{(1)}$, we have background fields ${\mathfrak{b}}_{I}$. Denote the partition function on 2+1d spacetime manifold ${M^{(3)}}$ with background field ${\mathfrak{b}}_{I}$ as ${\cal Z}_{\cal T}[{M^{(3)}},{\mathfrak{b}}_{I}]$. Then, the gauging corresponds to making ${\mathfrak{b}}_{I}$ dynamical by summing over all possible configurations of ${\mathfrak{b}}_{I}$:
@@
where $\cal G$ is the volume of the gauge group (see, e.g., App. B in [37]). Since ${\cal A}_{0}$ is non-anomalous (meaning that the theory coupled to ${\mathfrak{b}}_{I}$ is gauge-invariant), the procedure is well-defined and leads to a new TQFT ${\cal T}/{\cal A}_{0}$, as illustrated above by the example of the $\mathrm{U}(1)_{mn^{2}}$ theory.
@@
Now, suppose ${M^{(3)}}$ takes the form of $\Sigma\times S^{1}$ where $\Sigma$ is a closed surface, and $S^{1}$ is the Euclidean time circle. For simplicity, let us take $\Sigma=S^{2}$ so $\cal T$ has a unique state. Quantizing the theory, one obtains the state of ${\cal T}/{\cal A}_{0}$ as a coherent superposition of all possible insertion of (point-like) defects in ${\cal A}_{0}$ on the state of ${\cal T}$, [^15] If the TQFT is embedded in a non-topological UV theory, these defects can be the IR avatars of massive particles, or anyons. i.e., a proliferation of ${\cal A}_{0}$ defects.
This picture is also convenient for understanding the effect of gauging algebraically. The spectrum of lines in $\mathcal{T}/\mathcal{A}_{0}$ can be obtained through the following three-step procedure: [^16] This procedure was first discussed in the context of RCFT as an extension of the chiral algebra [73]. It was later related to a quotient of the gauge group of the corresponding Chern-Simons gauge theory [74]. (In the example leading to (3.25), we used a $\mathbb{Z}_{n}$ quotient of $\mathrm{U}(1)$.) From a more modern perspective, this is gauging a one-form symmetry [37]. The extension of the chiral algebra also leads to examples (e.g., ${\rm SU}(2)_{10}\to{\rm Spin}(5)_{1}$) that can be thought of as gauging a non-invertible symmetry. The term condensation for this procedure was introduced in [75] and is used often. However, it has led to some confusion. This issue will be summarized in [76], where the similarities and the distinctions between gauging and condensation will be clarified.
@@
3.3 TQFTs vs. gapped phases
@@
  • The traditional parameters affect correlations of local point operators at separated points. Examples include fermion masses, $\theta$-parameters, Yukawa coupling constants, etc.
@@ 4 From a response theory to a one-form symmetry
The response of the quantum Hall state to external electric and magnetic fields given by Eqns. (2.1), (2.2), and (2.3) can be compactly summarized by writing an effective theory for a background electromagnetic gauge field $A$:
@@
That this is the correct response theory can be seen by taking derivatives with respect to the spatial components $A_{i}$ to reproduce Eqns. (2.1), (2.2), or with respect to the time component $A_{0}$ to reproduce (2.3).
@@
The action (4.1) is well-defined for topologically trivial gauge fields. However, for generic $\sigma_{H}$, it is not meaningful when $A$ carries nonzero fluxes. Following [32], we can try to define it by viewing our 2+1d (arbitrary oriented) spacetime ${M^{(3)}}$ as the boundary of a 3+1d bulk ${M^{(4)}}$, and define the action (4.1) as
[…]
@@
Before we answer this question, we observe that the theory (4.2) is independent of the bulk for the following values of $\sigma_{H}$: for bosonic theories it must be an even integer, for fermionic/spin theories it can be any integer, and for spinc theories it can also be any integer. [^18] More precisely, this statement is true only if we add an appropriate term depending on the metric, which represents the thermal Hall conductivity. These values of $\sigma_{H}$ correspond exactly to the allowed ‘integer’ QHE states for bosons and fermions. Thus, it seems that the response theory of (4.1) is well-defined only for the integer states.
@@
Then (4.1) is not a meaningful 2+1d theory on ${M^{(3)}}$. This happened because we have integrated out the modes of a nontrivial TQFT. We refer to this TQFT as ${\cal T}$ [^19] Here we assume that our FQH state is gapped.
@@
[…]
@@
with $\lambda$ a properly normalized $\mathrm{U}(1)$ gauge field. (Here, we assume for simplicity that $pq$ is even. Otherwise, we should replace $q$ in (4.4) with $2q$.) This means that the only information in $A$ is the same as the information in a gauge field for a $\mathbb{Z}_{q}^{(1)}$ one-form symmetry. The corresponding $\mathbb{Z}_{q}^{(1)}$ gauge invariance is violated when ${M^{(4)}}$ has a boundary. Then, depending on the boundary conditions, this $\mathbb{Z}_{q}^{(1)}$ can act as a global symmetry. Even though this is not a symmetry of our problem, we will soon see that a related one-form symmetry must be present.
Motivated by this, we assume that the TQFT ${\cal T}$, has a $\mathbb{Z}_{s}^{(1)}$ global one-form symmetry for some $s$. (We will discuss the most general case in App. B.) As discussed in Sec. 3.2, in this case, it is natural to couple $\cal T$ to a background $\mathbb{Z}_{s}$-valued gauge field $\cal B$. In general, the one-form symmetry has an ’t Hooft anomaly, which can be characterized by an anomaly theory in one higher dimension. In our case, it is
@@
with ${\cal P}({\cal B})$ the Pontryagin square operation. (See Sec. 3.2.) As above, our 2+1d spacetime ${M^{(3)}}$ is the boundary of the 3+1d manifold ${M^{(4)}}$. [^20] Note that ${M^{(4)}}$ does not have to be physical; it is merely a mathematical device. Having said that, we can view the bulk ${M^{(4)}}$ as an SPT phase of a one-form $\mathbb{Z}_{s}^{(1)}$ symmetry, classified by an integer $r\text{ mod }2s$ with even $rs$. The corresponding boundary 2+1d theory has ’t Hooft anomaly for the $\mathbb{Z}_{s}^{(1)}$ one-form symmetry that is also labeled by the same integer $r\text{ mod }2s$. We emphasize that $r$ and $s$ are not necessarily co-prime.
So far, we have not yet related the expression (4.2) for $A$ and the expression (4.5) for $\cal B$. Before doing that, we remind the reader that as we said after (3.23), while (4.2) is well-defined even when ${M^{(4)}}$ has a boundary ${M^{(3)}}=\partial{M^{(4)}}$, this is not the case with (4.5). The “problem” with (4.5) associated with the boundary ${M^{(3)}}=\partial{M^{(4)}}$ is exactly canceled by the anomaly of the theory $\cal T$ on ${M^{(3)}}$.
As we mentioned around (4.4), the expression (4.2) has a one-form gauge symmetry, which is violated in the presence of a boundary. As we will soon see, this phenomenon is closely related to the anomaly in $\cal B$ and explains how (4.2) is related to it.
@@
Either way, we see that the two-form background field $\cal B$ is determined by the $\mathrm{U}(1)$ background gauge field $A$. (See e.g., [70, 80, 81, 38]).
@@
(For $\cal B$ given by (4.6), $\cal B$ has an integer lift and the Pontryagin square in (4.5) is simply a product of differential forms [72]. Compare with the discussion around (3.18)-(3.23).) This expression can be interpreted as the boundary term (4.1), provided
@@
[…]

### v3

@@ 3.1 Introduction to TQFTs
A basic physical operation is the fusion of two anyons. We denote the fusion of anyons $a$ and $b$ as $a\times b$. In general, $a\times b$ is a linear combination of anyons. If $a\times b$ is a single anyon, for all $b$, we refer to $a$ as an Abelian anyon. It is clear that the set of all Abelian anyons $\cal A$ forms an Abelian group whose multiplication is given by fusion. The $n$th power of an Abelian anyon $a^{n}$ is then defined in the obvious way.
@@
Let us return to the full set of anyons, not necessarily in $\cal A$. In a bosonic system, we require that $B$ be non-degenerate, i.e., the only anyon that braids trivially with all others is the identity $1$. (This requirement leads to the fact that the theory is modular; hence, the M in MTC.) A fermionic theory is not modular. It includes a unique anyon $c$, besides $1$, that braids trivially with all the other anyons. It is the physical fermion, which satisfies $c^{2}=1$ and $h(c)=\frac{1}{2}$. Often we identify $c$ with $1$, since both are local particles.
Let us demonstrate the general discussion with the special case of Abelian theories. They are described by a $K$-matrix Chern-Simons Lagrangian [^7] There is a procedure for mapping the algebraic data of the TQFT to a $K$ matrix, which is given in Appendix C of [64]. We will not review it here, except to mention that it is possible to go back and forth.
[…]
@@
Let us turn to the more general case. The topological spin $h(a)$ and the $\mathrm{U}(1)$ charge $Q(a)$ are subject to identifications because of attachment of local excitations. We consider three cases.
@@
    Physically, the ambiguity in the charge arises from attaching a transparent boson with integer charge.
@@
    In a spin theory where $c$ is identified with the vacuum, [^9] A spin TQFT is a TQFT that is well defined only on spin manifolds. And then, it depends on the choice of spin structure on that manifold [67, 68, 69]. Such a system does not have to be enriched by any global symmetry.
@@
    We prove in Appendix C that the different ways of assigning $\mathrm{U}(1)$ charge in this theory are essentially identical to the bosonic theory. They also correspond to different choices of Abelian vison, modulo fusion with $c$.
@@
    If $c$ is identified with the vacuum, then there is an additional equivalence relation which relates the tuple $(h(a),Q(a))$ to $(h(a\times c),Q(a\times c))$,
@@
    Just as $\hat{h}$, the $\pi$-flux defect is not just a convenient tool, but plays an important role in the physics. Classifying the different ways of assigning $\mathrm{U}(1)$ charges in a spinc TQFT can be done by extending the theory to a bosonic TQFT. Here, the $\pi$-flux line appears not as a defect line, but as an actual anyon in the theory. Once the theory has been extended to a bosonic TQFT, the techniques of [33] can be employed to classify charge assignments. We give much more detail in Appendix C.
@@ 3.2 Global symmetries, anomalies, and gauging
We now rephrase the discussion using the language of higher-form symmetries [39], within the framework of TQFT. The advantage of doing so is that the approach streamlines the matching of symmetries and anomalies between UV and IR theories, and the reasoning can be easily extended to nontopological theories with such symmetries.
@@
with $\lambda$ a flat $\mathrm{U}(1)$ gauge field, i.e., $d\lambda=0$, with $e^{is\oint\lambda}=1$ around every one-cycle. Such $\lambda$ is a $\mathbb{Z}_{s}$ gauge field and the global symmetry transformation (3.16) is a $\mathbb{Z}_{s}^{(1)}$ one-form global symmetry. [^12] Throughout this paper, we use the convention that $\mathbb{Z}_{s}^{(p)}$ denotes a $\mathbb{Z}_{s}$ $p$-form symmetry.
[…]
@@
A standard way to define the action based on (3.14) is to view our spacetime as the boundary of an auxiliary four manifold $M^{(4)}$} and to write (3.14) as ${s\over 4\pi}\int_{M^{(4)}}db\wedge db$ [32]. Indeed, we will frequently want to take advantage of this useful tool of viewing our 2+1d spacetime as the boundary, ${M^{(3)}}=\partial{M^{(4)}}$, of a 3+1d bulk, $M^{(4)}$}. We emphasize that the 3+1d manifold $M^{(4)}$} is not physical. It is merely a mathematical device.
[…]
@@
[…]
sdb=dAsdb=dA\, (3.21)
@@
Let us try to gauge the one-form symmetry. The anomaly prevents us from doing it. However, if there are $m\in 2\mathbb{Z}$ and $n\in\mathbb{Z}$ such that $s=mn^{2}$, then, we can gauge $\mathbb{Z}_{n}^{(1)}\subset\mathbb{Z}_{s}^{(1)}$, because the anomaly for this subgroup vanishes. Denote the gauge field for this subgroup as $\mathfrak{b}$, in terms of which ${\cal B}={s\over n}\mathfrak{b}$. [^14] We use uppercase characters for classical background fields and lowercase characters for dynamical fields. Clearly, $\mathfrak{b}$ is a $\mathbb{Z}_{n}$ gauge field. Then, for fixed $\mathfrak{b}$, (3.17) becomes
@@
For example, the $\mathrm{U}(1)_{8}$ TQFT has a $\mathbb{Z}_{8}^{(1)}$ one-form global symmetry. A $\mathbb{Z}_{2}^{(1)}$ subgroup of it is anomaly free and can be gauged. In the notation above, this corresponds to $s=8$ with $m=n=2$. This gauging leads to the $\mathrm{U}(1)_{2}$ TQFT.
We now generalize the gauging procedure and continue to focus on bosonic theories. A subset of the anyons in the TQFT $\cal T$ is the Abelian anyons ${\cal A}\subset{\cal T}$. They are the symmetry lines of the one-form global symmetry of $\cal T$. We consider an anomaly free subset of them ${\cal A}_{0}\subset{\cal A}$. This means that for all $x,y\in{\cal A}_{0}$,
@@
Gauging ${\cal A}_{0}$ leads to another TQFT denoted by ${\cal T}/{\cal A}_{0}$. To find it, we couple $\cal T$ to background gauge fields for ${\cal A}_{0}$. If the symmetry associated with ${\cal A}_{0}$ is $\otimes_{I}\mathbb{Z}_{n_{I}}^{(1)}$, we have background fields ${\mathfrak{b}}_{I}$. Denote the partition function on 2+1d spacetime manifold $M^{(3)}$} with background field ${\mathfrak{b}}_{I}$ as ${\cal Z}_{\cal T}[{M^{(3)}},{\mathfrak{b}}_{I}]$. Then, the gauging corresponds to making ${\mathfrak{b}}_{I}$ dynamical by summing over all possible configurations of ${\mathfrak{b}}_{I}$,
@@
where $\cal G$ is the volume of the gauge group (see, e.g., Appendix B in [39]). Since ${\cal A}_{0}$ is non-anomalous (meaning that the theory coupled to ${\mathfrak{b}}_{I}$ is gauge invariant), the procedure is well defined and leads to a new TQFT ${\cal T}/{\cal A}_{0}$, as illustrated above by the example of the $\mathrm{U}(1)_{mn^{2}}$ theory.
@@
Now, suppose $M^{(3)}$} takes the form of $\Sigma\times S^{1}$ where $\Sigma$ is a closed surface, and $S^{1}$ is the Euclidean time circle. For simplicity, let us take $\Sigma=S^{2}$ so $\cal T$ has a unique state. Quantizing the theory, one obtains the state of ${\cal T}/{\cal A}_{0}$ as a coherent superposition of all possible insertion of (point-like) defects in ${\cal A}_{0}$ on the state of $\cal T$}, [^15] If the TQFT is embedded in a nontopological UV theory, these defects can be the IR avatars of massive particles, or anyons. i.e., a proliferation of ${\cal A}_{0}$ defects.
This picture is also convenient for understanding the effect of gauging algebraically. The spectrum of lines in $\mathcal{T}/\mathcal{A}_{0}$ can be obtained through the following three-step procedure: [^16] This procedure was first discussed in the context of RCFT as an extension of the chiral algebra [75]. It was later related to a quotient of the gauge group of the corresponding Chern-Simons gauge theory [76]. [In the example leading to (3.25), we used a $\mathbb{Z}_{n}$ quotient of $\mathrm{U}(1)$.] From a more modern perspective, this is gauging a one-form symmetry [39]. The extension of the chiral algebra also leads to examples [e.g., ${\rm SU}(2)_{10}\to{\rm Spin}(5)_{1}$] that can be thought of as gauging a non-invertible symmetry. The term condensation for this procedure was introduced in [77] and is used often. However, it has led to some confusion. This issue will be summarized in [78], where the similarities and the distinctions between gauging and condensation will be clarified.
@@
3.3 TQFTs vs gapped phases
@@
  • The traditional parameters affect correlations of local point operators at separated points. Examples include fermion masses, $\theta$ parameters, Yukawa coupling constants, etc.
@@ 4 From a response theory to a one-form symmetry
The response of the quantum Hall state to external electric and magnetic fields given by Eqs. (2.1)–(2.3) can be compactly summarized by writing an effective theory for a background electromagnetic gauge field $A$,
@@
That this is the correct response theory can be seen by taking derivatives with respect to the spatial components $A_{i}$ to reproduce Eqs. (2.1) and (2.2), or with respect to the time component $A_{0}$ to reproduce (2.3).
@@
The action (4.1) is well defined for topologically trivial gauge fields. However, for generic $\sigma_{H}$, it is not meaningful when $A$ carries nonzero fluxes. Following [32], we can try to define it by viewing our 2+1d (arbitrary oriented) spacetime $M^{(3)}$} as the boundary of a 3+1d bulk $M^{(4)}$}, and define the action (4.1) as
${\sigma_{H}\over 4\pi}\int_{M^{(4)}}dA\wedge dA\,.$ (4.2)
@@
Before we answer this question, we observe that the theory (4.2) is independent of the bulk for the following values of $\sigma_{H}$: for bosonic theories it must be an even integer, for fermionic/spin theories it can be any integer, and for spinc theories it can also be any integer. [^18] More precisely, this statement is true only if we add an appropriate term depending on the metric, which represents the thermal Hall conductivity. These values of $\sigma_{H}$ correspond exactly to the allowed ‘integer’ QHE states for bosons and fermions. Thus, it seems that the response theory of (4.1) is well defined only for the integer states.
@@
Then (4.1) is not a meaningful 2+1d theory on $M^{(3)}$}. This happened because we have integrated out the modes of a nontrivial TQFT. We refer to this TQFT as $\cal T$}. [^19] Here we assume that our FQH state is gapped.
@@
We pause to note that the discussion of flux threading in Sec. 2.2 is a manifestation of the fact that (4.1) is globally ill defined. Consider a one-parameter family of background gauge fields depending on a parameter, Lorentzian time, or Euclidean time, shifting $\oint A_{\varphi}d\varphi$ by $2\pi$. This is a closed path in the space of $A$’s, but the integral of (4.1) is not single valued under it. Instead, it is shifted by ${\sigma_{H}\over 2\pi}\int d\varphi drdtE_{r}$, signaling the transfer of charge $\sigma_{H}$ from one side of the annulus to the other.
@@
with $\lambda$ a properly normalized $\mathrm{U}(1)$ gauge field. [Here, we assume for simplicity that $pq$ is even. Otherwise, we should replace $q$ in (4.4) with $2q$.] This means that the only information in $A$ is the same as the information in a gauge field for a $\mathbb{Z}_{q}^{(1)}$ one-form symmetry. The corresponding $\mathbb{Z}_{q}^{(1)}$ gauge invariance is violated when $M^{(4)}$} has a boundary. Then, depending on the boundary conditions, this $\mathbb{Z}_{q}^{(1)}$ can act as a global symmetry. Even though this is not a symmetry of our problem, we will soon see that a related one-form symmetry must be present.
[…]
@@
with ${\cal P}({\cal B})$ the Pontryagin square operation. (See Sec. 3.2.) As above, our 2+1d spacetime $M^{(3)}$} is the boundary of the 3+1d manifold $M^{(4)}$}. [^20] Note that $M^{(4)}$} does not have to be physical; it is merely a mathematical device. Having said that, we can view the bulk $M^{(4)}$} as an SPT phase of a one-form $\mathbb{Z}_{s}^{(1)}$ symmetry, classified by an integer $r\text{ mod }2s$ with even $rs$. The corresponding boundary 2+1d theory has ’t Hooft anomaly for the $\mathbb{Z}_{s}^{(1)}$ one-form symmetry that is also labeled by the same integer $r\text{ mod }2s$. We emphasize that $r$ and $s$ are not necessarily coprime.
[…]
@@
[…]
@@
We conclude that the response theory (4.1) is reproduced by a TQFT with a $\mathbb{Z}_{s}^{(1)}$ one-form symmetry and anomaly labeled by $r$ such that $r/s=p/q\text{ mod }2$. [^21] A related discussion of the fractional part of the coefficient of a Chern-Simons term as a diagnostic of a nontrivial IR theory has appeared (in the context of duality and gapless theories) in [81]. In App. B we will show that what we have done is essentially the only way to reproduce (4.1).